### v1

# Leggett-Garg Inequality for Bosons in a Double-Well Potential

Abstract
The Leggett-Garg inequality (LGI) serves as a criterion to determine the adherence of macroscopic system dynamics to macrorealism, as postulated by Leggett and Garg. A violation of this inequality implies either the absence of a realistic description of the system or the impracticality of noninvasive measurements. In this work, we investigate the violation of the LGI for the system of bosons in a double-well potential. Specifically, we explore the violation of the LGI in the dynamics of bosons in a double-well potential in the Bose-Einstein-Condensation (BEC) regime, where the system can be considered as two weakly coupled Bose condensates, and in the single-particle regime to establish the conditions under which the violation of the LGI occurs. Our analysis reveals that the LGI is violated due to Josephson oscillations, while it remains unviolated in the strong coupling regime, attributed to the self-trapping phenomena. Notably, we observe that the violation of the LGI becomes increasingly significant as the particle number increases. These findings provide valuable insights into the macrorealistic behavior of Bose condensates and highlight the effect of measurements on the dynamics of a macroscopic system.
Introduction.—How classical behaviors of objects, which obey our intuition about how the macroscopic world behaves, can be distinguished from non-classical behaviors such as what are described by quantum theory? The Leggett-Garg inequality (LGI) was proposed to answer this question [1, 2, 3]. It serves as a test of a concept called macrorealism, which consists of two conditions: macrorealism per se (MRPS) and noninvasive measurability (NIM). MRPS assumes that a macroscopic system with two or more macroscopically distinct states available to it is always in one or the other of these states, and NIM assumes that it is possible to determine the state of the system with arbitrary small perturbations on its subsequent dynamics. A violation of the LGI implies a departure from macrorealism. The exploration of the LGI was inspired by the question of whether macroscopic coherence, as illustrated by Schrödinger’s cat gedanken experiment, can be realized in a laboratory [4].
The violation of the LGI has been studied in several systems, experimentally and theoretically [3, 7, 5, 6, 8, 9, 10, 11, 12, 13, 14, 15]. Recently, experimental LG tests have been achieved in a superconducting flux qubit with genuine macroscopicity, refuting its classical realistic description [16].
The system of Bose-Einstein condensates (BECs) in a double-well potential is suitable for testing the LGI to clarify the boundary between classical and quantum regimes. The dynamics of a BEC in a double-well potential has been observed in several experiments [19, 21, 20, 22, 23, 17, 18, 25, 24]. In particular, in the experiment of the Heidelberg group [19], Josephson oscillations and macroscopic quantum self-trapping were successfully observed, and the experimental result exhibits agreement with the prediction by the Gross-Pitaevskii (GP) equation [28, 27, 26]. It implies that a population imbalance and a phase difference between the two wells have definite values independent of measurements in the presence of sufficiently large number of particles in the double-well potential [29].
In this letter, we discuss the violation of the LGI for bosons in a double-well potential. We find the violation of the LGI in the BEC regime, in which the system can be considered as two weakly coupled BECs, despite the fact that their dynamics are well described by coupled semiclassical equations for definite phase and population differences, which are presumed to be unaffected by measurements. We reveal that the violation of the LGI results from the collapse of the state due to projective measurements in the intermediate time, which necessarily affects the subsequent dynamics. We find that the violation of the LGI becomes increasingly significant as the particle number increases. Furthermore, we find that the critical value of interaction strength for the onset of the violation of the LGI coincides with the one for self-trapping. In the single-particle regime, in which bosons do not interact each other, the bosons tunneling backwards induced by the measurement are found to be crucial for the violation of the LGI.
Model.—The system of bosons in a double-well potential is well described by the Bose-Josephson junction (BJJ) model [28, 33, 31, 35, 34, 37, 41, 36, 38, 30, 39, 40, 32]. Using the Schwinger boson representation $\hat{S}_{x}=(\hat{a}_{\rm L}^{\dagger}\hat{a}_{\rm R}+\hat{a}_{\rm R}^{\dagger}\hat{a}_{\rm L})/2$ and $\hat{S}_{z}=(\hat{n}_{\rm L}-\hat{n}_{\rm R})/2$, where $\hat{a}_{\alpha}$ ($\hat{a}_{\alpha}^{\dagger}$) and $\hat{n}_{\alpha}=\hat{a}^{\dagger}_{\alpha}\hat{a}_{\alpha}$ are annihilation (creation) and number operators of a boson in the well $\alpha={\rm L},{\rm R}$, the Hamiltonian of the BJJ model can be written as [36, 38, 30]
$\hat{H}_{\rm BJJ}=-J\hat{S}_{x}+\frac{U}{N}\hat{S}_{z}^{2},$ (1)
where $J$ is the hopping strength, $U$ is on-site interaction and $N$ is the total number of particles. Hereafter, we set $\Lambda=U/J$ and $\hbar=1$.
Method.—The LGI is formulated for the correlation functions $C_{ij}=\langle Q_{i}Q_{j}\rangle$ of dichotomic variables $Q_{i}=\pm 1$, which is given as the result of the measurement at time $t_{i}$, where $i,j=1,2,3$. MRPS ensures the existence of the value of the observable $Q_{i}$ at $t_{i}$, regardless of whether the measurement is performed or not. Thus we can define the joint probabilities $P_{ij}(Q_{1},Q_{2},Q_{3})$, where measurements are performed at $t_{i}$ and $t_{j}$, whereas the system is unmeasured at time $t_{k}$ ($\neq t_{i},t_{j}$). Under only MRPS, the joint probability may depend on when measurements are made, since measurements at different times can affect the time-evolution differently. By adding the NIM condition, however, the subscripts $i$ and $j$ can be dropped since any measurement does not affect the system under NIM. Then the joint probability $P(Q_{1},Q_{2},Q_{3})$ yields all correlation functions and we obtain the LGI [1, 3]
$LG\equiv C_{12}+C_{23}-C_{13}\leq 1.$ (2)
We follow the dynamics of the system numerically by exact diagonalization of the Hamiltonian (1). In evaluation of the LGI, in this work, we define the dichotomic observable as $\hat{Q}={\rm sgn}(\hat{S}_{z})$, where ${\rm sgn}(x)=1$ if $x\geq 0$, otherwise ${\rm sgn}(x)=-1$. For the sake of simplicity, we set an initial state at the time $t_{1}$. Throughout this letter, we focus on the identical time interval and define $\tau=t_{2}-t_{1}=t_{3}-t_{2}$. Then all three two-time correlation functions depend only on the time interval $\tau$. Based on the above setting, we calculate $LG$ defined as $LG=C_{12}+C_{23}-C_{13}$ (also see the Supplemental Material (SM) for details on the calculations of the correlation functions).
BEC regime.—First, we examine violation of the LGI in the BEC regime $(\Lambda\neq 0$ and $N\gg 1)$, in which the system can be considered as two weakly coupled Bose condensates. Such Bose condensates have been studied by a semiclassical approach based on two coupled GP equations, which predict Josephson oscillation or self-trapping depending on the value of $\Lambda$ and population imbalance at initial time [19, 21, 28, 30, 43, 42].
[…]
Figure 1: (a)-(e) Time evolutions of the occupation probabilities of the left well $P(n_{\rm L})$ for $\Lambda=0.0$ (a), $1$ (b), $1.5$ (c), $1.9$ (d), and $2.5$ (e). Here we set $N=2000$. (f)-(h) Colormaps of $LG$ for $N=100$ (f), $200$ (g) and $2000$ (h). The two black dashed lines in (f)-(h) denote the critical value of self-trapping $|\Lambda_{\rm c}|=2$. (i) Violation region of $\Lambda$. The light-blue region denotes the violation region, while the gray region non-violation region. The red dashed line denotes the boundary between them, which coincides with the onset of self-trapping.
Population dynamics in the BEC regime.—Figs. 1-1 present the time-evolutions of the occupation probability of the left well $P(n_{\rm L})$ in the case of $N=2000$ with the initial state $|N\rangle_{\rm L}|0\rangle_{\rm R}$. The peak of the occupation probability oscillates between $n_{\rm L}=0$ and $n_{\rm L}=N$ due to single-particle tunneling effect when $\Lambda=0$ (Fig. 1). A similar coherent oscillation occurs in the presence of on-site interaction, but it is rather induced by Josephson effect when $20<\Lambda<$ 2 (Figs. 1-1). For strong interaction strength, the Josephson oscillation ceases and self-trapping occurs when $\Lambda\geq 2$, where the peak of the occupation probability oscillates keeping the population imbalance (Fig. 1). The critical value of interaction strength for the onset of self-trapping $\Lambda_{\rm c}=2$ agrees with the value derived in the semiclassical analysis based on the GP equation[28].
If the dynamics obeys coupled GP equations, oscillation of the population imbalance continues without damping [28, 31, 42]. The damping of the oscillation that occurs in the presence of on-site interaction in Figs. 1-1 can be understood as being due to the intrinsic quantum fluctuation in the initial state [30]. That is, since the initial population imbalance is fixed, the initial phase difference between the Bose condensates should be fluctuating due to the uncertainty relation between population imbalance and phase difference [30]. As $n_{\rm L}$ and $n_{\rm R}$ becomes larger, the quantum fluctuation ($\simeq 1/\sqrt{n_{{\rm L},{\rm R}}}$ [44]) becomes more negligible [38, 28]. This is the reason why, with the increase of $N$, the coherent Josephson oscillation persists for longer (see Fig. S3 in the SM).
BECs in a double-well potential are considered to possess macroscopic observables for sufficiently large $N$. In fact, the damping is suppressed for larger $N$, and the population dynamics is consistent with the semiclassical description. The experimental data in Ref. [19] have also been shown to be consistent with the prediction of the semiclassical analysis. Therefore, both the theoretical and experimental studies support the existence of macroscopic wave functions of two BECs with definite phase difference and population imbalance for sufficiently large $N$. Upon such considerations, it seems reasonable to expect non-violation of the LGI.
Violation of the LGI in the BEC regime.—Fig. 1 shows the result of the LG test in the BEC regime for $N=2000$: $LG$ is plotted as a function of $\tau$ and $\Lambda$. $LG$ exhibits periodic oscillations as a function of $\tau$ that reflects the oscillation of the occupation probabilities in Figs. 1-1. Contrary to the expectation, the LGI is violated near the peaks of the oscillations for the interaction $-2<\Lambda<2$, where Josephson oscillations take place for $\Lambda\neq 0$, while single-particle oscillations for $\Lambda=0$. It implies that macrorealism is excluded even for such a large $N$, and thus contradicts what is expected by the GP description; the LGI is never broken for the GP description [29].
[…]
Figure 2: Effect of collapse of the state due to measurement at time $t_{2}$. (a) Plots of $LG$ in Eq. (2) as a function of $\tau J$. The upper blue region represents a violation of the LGI in Eq. (2). The blue (red) line denotes the plot of $LG$ where the state is (not) collapsed at $t_{2}$. (b) $|\langle Q_{3}\rangle_{{\rm M}_{2}}-\langle Q_{3}\rangle|$ as a function of $\tau J$. The parameters used in (a) and (b) are $N=100$ and $\Lambda=1$.
As shown in Fig. 1, when $\Lambda$ exceeds the critical value indicated by the dashed line, the LGI becomes non-violated regardless of $\tau$. It is noteworthy that this critical value coincides with the one for the onset of self-trapping $\Lambda_{\rm c}=2$ (see Fig. 1). The self-trapping prevents the system from coherent oscillation and yields, $C_{12}=C_{13}=1$ for any interval $\tau$ (see SM). Then $C_{12}$ and $C_{13}$ cancel out and $LG=C_{23}\leq 1$. Thus, the LGI is non-violated when self-trapping occurs.
Fig. 1 shows that the period of oscillation of $LG$ increases and its amplitude decreases as $\Lambda\to\Lambda_{\rm c}^{-}$. It is due to the softening of the Josephson oscillation associated with the transition from the Josephson regime to the self-trapping regime [38, 45].
The LGI becomes non-violated for $\tau$ greater than a characteristic value when $0<\Lambda<\Lambda_{\rm c}$, which becomes smaller as the interaction strength increases but it becomes longer as $N$ increases, as shown in Figs. 1-1. This behavior is due to the damping of Josephson oscillation, which destroys the correlation in time. The Josephson oscillation indeed decays after a characteristic time, which becomes smaller as the interaction strength increases, as shown in Figs. 1-1, but it becomes longer as $N$ increases due to suppression of fluctuation of phase difference of two Bose condensates.
The LGI is found to be violated symmetrically between repulsive ($\Lambda>0$) and attractive ($\Lambda<0$) on-site interaction in Figs. 1-1 since the Hamiltonian with repulsive interaction can be mapped to the one with attractive interaction as $\hat{R}_{z}(\pi)\hat{H}_{\rm BJJ}(\Lambda)\hat{R}_{z}^{\dagger}(\pi)=-\hat{H}_{\rm BJJ}(-\Lambda)$, where $\hat{R}_{z}(\theta)=\exp\quantity(-i\theta\hat{S}_{z})$. From this relationship, we obtain
$P(n_{\rm L},\Lambda,t)=P(n_{\rm L},-\Lambda,t),$ (3)
where $P(n_{\rm L},\Lambda,t)$ denotes the probability of $n_{\rm L}$ bosons on the left-well at time $t$ for $\Lambda$ (see SM). This symmetric relationship between repulsive and attractive on-site interaction results in the symmetric evolution of $LG$ between them in Figs. 1-1.
Origin of the violation of the LGI.—The violation of the LGI in Fig. 1 results from the collapse of the state due to projective measurements made at time $t_{2}$. As shown in Fig. 2, the LGI holds if the state is assumed not to be collapsed by measurements at $t_{2}$, while if the state is collapsed at $t_{2}$ the LGI is violated. The effect of the collapse appears explicitly in Fig. 2 in the comparison between $\langle Q_{3}\rangle_{{\rm M}_{2}}$, where the state is assumed to be collapsed due to measurements at $t_{2}$, and $\langle Q_{3}\rangle$, where measurements are not made at $t_{2}$. The condition $|\langle Q_{3}\rangle_{{\rm M}_{2}}-\langle Q_{3}\rangle|\neq 0$ implies that the dynamics after time $t_{2}$ evolves differently depending on whether a measurement is made at $t_{2}$ or not. Comparison between Figs. 2 and 2 shows that the LGI is violated for $\tau$ at which $|\langle Q_{3}\rangle_{{\rm M}_{2}}-\langle Q_{3}\rangle|$ becomes large. Therefore $Q_{3}$ of $C_{13}$ and that of $C_{23}$ evolve differently, and this fact leads to violations of the LGI. As shown in Figs. 2 and 2, the effect of collapse of the state on $LG$ and $\langle Q_{3}\rangle$ due to measurements at $t_{2}$ disappears as the oscillation of population decays.
[…]
Figure 3: (a) Time evolutions of $P(n_{\rm L})$ for $z(0)=0.28$. (b) Violation of the LGI for $z(0)=0.2$ (the red line), 0.28 (the blue line), and 0.4 (the green line). These parameters are well below the critical value of macroscopic quantum self-trapping. In (a) and (b), the parameters are $N=1000$ and $\Lambda=15$.
Observability of violation of the LGI.—We numerically test the LGI for the nearly same parameters for which Josephson oscillations were observed in Ref. [19]. Here, we set $\Lambda=15$ and $N=1000$. The LGI is shown to be broken for realistic parameters in Fig. 3, in which the evolution of $LG$ is plotted as a function of $\tau$ for several initial population imbalance $z(0)=(n_{\rm L}(0)-n_{\rm R}(0))/N$. $z(0)$ needs to be smaller than the critical value for self-trapping $z_{\rm c}(0)\simeq 0.5$ [31]. $z(0)=0.28$ is expected to be more desirable for the demonstration of violation of the LGI. Fig. 3 shows the time evolution of $P(n_{\rm L})$ for $z(0)=0.28$. It exhibits tunneling of atoms between the two wells, which indeed leads to the violation of the LGI.
[…]
Figure 4: (a) Violation of the LGI. (b) Three correlation functions $C_{12}$ (the red dashed line), $C_{23}$ (the black solid line) and $C_{13}$ (the green dotted line) as functions of $\tau$. (c)/(d) Time evolution of $P(n_{\rm L})$ in the absence/presence of the measurement at $tJ/2\pi=0.2$ (represented by the gray dotted line). $n_{\rm L}=70$ at $tJ/2\pi=0.2$ in (d). (e) Colormap of $LG$ for all different initial occupation number in the left well $n_{\rm L}(0)$. The violation of the LGI is symmetric under the reflection with respect to the line $n_{\rm L}(0)=50$. The blue color in (e) shows the part where the LGI is violated. In (a)-(e), the parameters are $N=100$ and $\Lambda=0$.
Single particle regime.—Next, we study violation of the LGI in the absence of on-site interaction. In this case the Hamiltonian is simply given by $\hat{H}_{\rm BJJ}=-J\hat{S}_{x}$. For this model, the Schrödinger equation is analytically solvable and the resultant wave function at time $t$ evolved from the initial state $|N-l\rangle_{\rm L}|l\rangle_{\rm R}$ is given by
|$\ket{\Psi(t)}=$ $\frac{1}{\sqrt{(N-l)!}}\frac{1}{\sqrt{l!}}\left(\hat{a}^{\dagger}_{\rm L}\cos\frac{J}{2}t-i\hat{a}^{\dagger}_{\rm R}\sin\frac{J}{2}t\right)^{N-l}$
×$\times\left(\hat{a}^{\dagger}_{\rm R}\cos\frac{J}{2}t-i\hat{a}^{\dagger}_{\rm L}\sin\frac{J}{2}t\right)^{l}|\mathrm{vac}\rangle.$ (4)
The first/second line of Eq. (Leggett-Garg Inequality for Bosons in a Double-Well Potential) represents the coherent oscillation of the bosons initially located in the left/right well. These bosons flow in opposite directions. The former represents bosons initially tunneling from the left well to the right well, as shown in Fig. 4, while the latter represents those initially tunneling from the right well to the left well. As shown in Fig. 4, the peak of the occupation probability coherently oscillates between $n_{\rm L}=0$ and $n_{\rm L}=N$ without damping. Due to this coherent oscillation the LGI is violated, as shown in Fig. 4.
Focusing on the first violation in Fig. 4, this behavior can be explained as follows: Until $C_{13}$ starts to decrease, the LGI holds, since $C_{12}=C_{13}=C_{23}=1$. The LGI begins to be first violated when $C_{13}$ starts decreasing at $\tau J/2\pi\simeq 0.125$. Since $C_{13}$ abruptly becomes $-1$ whilst $C_{12}=1$ and $C_{23}$ is still positive, the LGI becomes violated quite abruptly. The value of the peak of $LG$ becomes larger with the increase of $N$, due to more abrupt decrease of $C_{13}$ (see SM).
The persistence of this violation is attributed to the bosons tunneling backwards in Fig. 4 induced by the measurement, which corresponds to the second line of Eq. (Leggett-Garg Inequality for Bosons in a Double-Well Potential). In the case of $C_{23}$ the bosons tunneling backwards impede the decrease of $C_{23}$, whilst such effect is absent in the case of $C_{13}$, where no measurement is performed at $t_{2}$ (see Fig. 4). The LGI again begins to hold approximately when $C_{12}$ becomes zero at $\tau J/2\pi\simeq 0.25$ (see Fig. 4).
We also test the LGI for all different initial occupation numbers of the left well $n_{\rm L}(0)$ with $N=100$. In Fig. 4, it can be seen that the peak value of $LG$ increases with a larger initial population imbalance (see SM).
Conclusion and discussion.—In this work, we study violations of the LGI for bosons in a double-well potential. We find that the LGI is violated due to Josephson oscillations in the weak coupling regime, while it remains unviolated in the strong coupling regime, attributed to the self-trapping phenomena. Remarkably, we observe that the LGI remains violated even in the large $N$ system. We also find that, in this system, on-site interaction induces a dephasing effect that destroys temporal correlations, resulting in the LGI no longer being violated afterwards. We have confirmed that the LGI is broken below the critical value of interaction strength for self-trapping. In contrast, it has been shown that the LGI is no longer violated when self-trapping takes place. Furthermore, we have revealed that the LGI is broken for nearly the same parameters for which the Josephson oscillations are observed in the experimental work of Ref. [19], thus indicating the presence of macroscopic quantum coherence in the semiclassical regime.
Our result suggests that either MRPS or NIM is violated, or possibly both, in the system of bosons in a double-well potential with a large number of particles. In our LG tests, we perform projection measurements, which necessarily affect the subsequent dynamics. Therefore, if measurements can be made without collapsing the state in any way [46], the result can be different from ours.
Acknowledgments.—We thank G. Kimura, M. Kunimi and T. Nikuni for fruitful discussions. ST thanks F. Dalfovo, G. Ferrari, and A. Recati for useful discussions. RY is partially supported by a Grant-in-Aid of MEXT for Scientific Research (KAKENHI Grant Nos. 19K14616 and 20H01838). ST is supported by the Japan Society for the Promotion of Science Grant-in-Aid for Scientific Research (KAKENHI Grant No. 19K03691).