### v5

# Violation of the Leggett-Garg inequality for dynamics of a Bose-Einstein condensate in a double-well potential

Abstract
The Leggett-Garg inequality (LGI) serves as a criterion to determine the adherence of macroscopic system dynamics to macrorealism, as introduced by Leggett and Garg. A violation of this inequality implies either the absence of a realistic description of the system or the impossibility of noninvasive measurement. In this Letter, we investigate the violation of the LGI for the system of bosons in a double-well potential. Specifically, we explore the violation of the LGI in the dynamics of bosons in a double-well potential in the Bose-Einstein-condensation (BEC) regime, where the system can be considered as two weakly coupled Bose condensates, and in the single-particle regime to establish the conditions under which the violation of the LGI occurs. Our analysis reveals that the LGI is violated due to Josephson oscillations, while it remains unviolated in the strong coupling regime, attributed to the self-trapping phenomena. Notably, we observe that the violation of the LGI becomes increasingly significant as the particle number increases. These findings provide valuable insights into the macrorealistic behavior of Bose condensates and highlight the effect of measurements on the dynamics of a macroscopic system.
Introduction. How can classical behaviors of objects, which obey our intuition about how the macroscopic world behaves, be distinguished from non-classical behaviors such as what are described by quantum theory? The Leggett-Garg inequality (LGI) was proposed to answer this question [1, 2, 3]. It serves as a test of a concept called macrorealism, which consists of two conditions: macrorealism per se (MRPS) and noninvasive measurability (NIM). MRPS assumes that a macroscopic system with two or more macroscopically distinct states available to it is always in one or the other of these states, and NIM assumes that it is possible to determine the state of the system with arbitrary small perturbations on its subsequent dynamics. A violation of the LGI implies a departure from macrorealism. The exploration of the LGI was inspired by the question of whether macroscopic coherence, as illustrated by Schrödinger’s cat gedanken experiment, can be realized in a laboratory [4].
The violation of the LGI has been studied in several systems, experimentally and theoretically [3, 7, 5, 6, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Recently, experimental LG tests have been achieved in a superconducting flux qubit with genuine macroscopicity, refuting its classical realistic description [18].
The system of Bose-Einstein condensates (BECs) in a double-well potential is suitable for testing the LGI to clarify the boundary between classical and quantum regimes. The dynamics of a BEC in a double-well potential has been observed in several experiments [21, 23, 22, 24, 25, 19, 20, 27, 26]. In particular, in the experiment of the Heidelberg group [21], Josephson oscillations and macroscopic quantum self-trapping were successfully observed, and the experimental result exhibits agreement with the prediction by the Gross-Pitaevskii (GP) equation [30, 29, 28]. It implies that a population imbalance and a phase difference between the two wells have definite values independent of measurements in the presence of sufficiently large number of particles in the double-well potential [31].
In this Letter, we discuss the violation of the LGI for bosons in a double-well potential. We find the violation of the LGI in the BEC regime, in which the system can be considered as two weakly coupled BECs, despite the fact that their dynamics are well described by coupled semiclassical equations for definite phase and population differences, which are presumed to be unaffected by measurements. We reveal that the violation of the LGI is associated with the collapse of the state due to projective measurements in the intermediate time, which necessarily affects the subsequent dynamics. We find that the critical value of interaction strength for the onset of the violation of the LGI coincides with the one for self-trapping. Furthermore, we observe that the violation of the LGI becomes increasingly significant as the particle number increases. In the single-particle regime, in which bosons do not interact with each other, the bosons tunneling backwards induced by the measurement are found to be crucial for the violation of the LGI.
[…]
Figure 1: (a)–(e) Time evolutions of the occupation probabilities of the left well $P(n_{\rm L})$ for $\Lambda=0$ (a), $1$ (b), $1.5$ (c), $1.9$ (d), and $2.5$ (e). Here, we set $N=2000$. (f) The light-blue/gray region denotes the violation/nonviolation region. The red dashed line denotes the critical value of the LG violation for $N=2000$. (g)–(i) Color maps of $K$ for $N=100$ (g), $500$ (h), and $2000$ (i). The two red dashed lines in each panel denote the critical values ±$\pm\Lambda_{\rm c}(N)$. (j), (k) Maximal violations of the LGI as functions of $\Lambda$ and $N$, respectively. In (j), the red, blue, and green lines correspond to $N=100$, $500$, and $2000$, respectively. The vertical black dashed line in (j) indicates $\Lambda_{\rm st}^{\rm GP}$. In (k), the blue and green dots correspond to $\Lambda=0$ and $1$, respectively.
Model. The system of bosons in a double-well potential is well described by the Bose-Josephson junction (BJJ) model [30, 35, 33, 37, 36, 39, 42, 38, 40, 32, 41, 34]. Using the Schwinger boson representation $\hat{S}_{x}=(\hat{a}_{\rm L}^{\dagger}\hat{a}_{\rm R}+\hat{a}_{\rm R}^{\dagger}\hat{a}_{\rm L})/2$ and $\hat{S}_{z}=(\hat{n}_{\rm L}-\hat{n}_{\rm R})/2$, where $\hat{a}_{\alpha}$ ($\hat{a}_{\alpha}^{\dagger}$) and $\hat{n}_{\alpha}=\hat{a}^{\dagger}_{\alpha}\hat{a}_{\alpha}$ are annihilation (creation) and number operators of a boson in the well $\alpha={\rm L},{\rm R}$, the Hamiltonian of the BJJ model can be written as [38, 40, 32]
$\hat{H}_{\rm BJJ}=-J\hat{S}_{x}+\frac{U}{N}\hat{S}_{z}^{2},$ (1)
where $J$ is the hopping strength, $U$ is the on-site interaction, and $N$ is the total number of particles. Hereafter, we set $\Lambda=U/J$ and $\hbar=1$.
Method.—The LGI is formulated for the correlation functions $C_{ij}=\langle Q_{i}Q_{j}\rangle$ of dichotomic variables $Q_{i}=\pm 1$, which is given as the result of the measurement at time $t_{i}$, where $i,j=1,2,3$. MRPS ensures the existence of the value of the observable $Q_{i}$ at $t_{i}$, regardless of whether or not the measurement is performed. Thus we can define the joint probabilities $P_{ij}(Q_{1},Q_{2},Q_{3})$, where measurements are performed at $t_{i}$ and $t_{j}$, whereas the system is unmeasured at time $t_{k}$ ($\neq t_{i},t_{j}$). Under only MRPS, the joint probability may depend on when measurements are made, since measurements at different times can affect the time-evolution differently. By adding the NIM condition, however, the subscripts $i$ and $j$ can be dropped since any measurement does not affect the system under NIM. Then the joint probability $P(Q_{1},Q_{2},Q_{3})$ yields all correlation functions and we obtain the LGI [1, 3].
$K\equiv C_{12}+C_{23}-C_{13}\leq 1.$ (2)
We follow the dynamics of the system numerically by exact diagonalization of the Hamiltonian (1). In evaluation of the LGI, in this work, we define the dichotomic observable as $\hat{Q}={\rm sgn}(\hat{S}_{z})$, where ${\rm sgn}(x)=1$ if $x\geq 0$, otherwise ${\rm sgn}(x)=-1$. For the sake of simplicity, we set an initial state at the time $t_{1}$. Throughout this Letter, we focus on the identical time interval and define $\tau=t_{2}-t_{1}=t_{3}-t_{2}$. Then all three two-time correlation functions depend only on the time interval $\tau$. Based on the above setting, we calculate $K$ in Eq. (2) [see also the Supplemental Material (SM) for details on the calculations of the correlation functions].
BEC regime. First, we examine violation of the LGI in the BEC regime ($\Lambda\neq 0$ and $N\gg 1$), in which the system can be considered as two weakly coupled BECs. Such BECs have been studied by a semiclassical approach based on two coupled GP equations, which predict Josephson oscillation or self-trapping depending on the value of $\Lambda$ and population imbalance at initial time [21, 23, 30, 32, 44, 43].
Population dynamics in the BEC regime. Figures 11 present the time-evolutions of the occupation probability of the left well $P(n_{\rm L})$ in the case of $N=2000$ with the initial state $|N\rangle_{\rm L}|0\rangle_{\rm R}$. The peak of the occupation probability oscillates between $n_{\rm L}=0$ and $n_{\rm L}=N$ due to the single-particle tunneling effect when $\Lambda=0$ [Fig. 1] A similar coherent oscillation occurs in the presence of an on-site interaction, but it is rather induced by the Josephson effect when $0<\Lambda<\Lambda_{\rm st}$ [Figs. 11]. For a strong interaction strength, the Josephson oscillation ceases and self-trapping occurs when $\Lambda>\Lambda_{\rm st}$, where the peak of the occupation probability oscillates keeping the population imbalance [Fig. 1]. The critical value of interaction strength for the onset of self-trapping $\Lambda_{\rm st}\simeq 2$ agrees well with the value derived in the semiclassical analysis based on the GP equation $\Lambda_{\rm st}^{\rm GP}=2$ [30].
If the dynamics obeys coupled GP equations, oscillation of the population imbalance continues without damping [30, 33, 43]. The damping of the oscillation that occurs in the presence of an on-site interaction in Figs. 11 can be understood as being due to the intrinsic quantum fluctuation in the initial state [32]. That is, since the initial population imbalance is fixed, the initial phase difference between the BECs should be fluctuating due to the uncertainty relation between the population imbalance and phase difference [32]. As $n_{\rm L}$ and $n_{\rm R}$ becomes larger, the quantum fluctuation ($\simeq 1/\sqrt{n_{{\rm L},{\rm R}}}$ [45]) becomes more negligible [40, 30]. This is the reason why, with the increase of $N$, the coherent Josephson oscillation persists for longer (see Fig. S3 in the SM).
BECs in a double-well potential are considered to possess macroscopic observables for sufficiently large $N$. In fact, the damping is suppressed for larger $N$, and the population dynamics is consistent with the semiclassical description (see SM). The experimental data in Ref. [21] have also been shown to be consistent with the prediction of the semiclassical analysis. Therefore, both the theoretical and experimental studies support the existence of macroscopic wave functions of two BECs with a definite phase difference and population imbalance for sufficiently large $N$. Upon such considerations, it seems reasonable to expect nonviolation of the LGI.
Violation of the LGI in the BEC regime. Figure 1 shows the result of the LG test in the BEC regime for $N=2000$: $K$ is plotted as a function of $\tau$ and $\Lambda$. $K$ exhibits periodic oscillations as a function of $\tau$ reflecting the oscillation of the occupation probabilities in Figs. 11. Contrary to the expectation, the LGI is violated near the peaks of the oscillations for the interaction $-\Lambda_{\rm c}<\Lambda<\Lambda_{\rm c}$. For $|\Lambda|\geq\Lambda_{\rm c}$ the LGI is nonviolated. Remarkably, the critical value $\Lambda_{\rm c}$ that separates violation and nonviolation of the LGI coincides with $\Lambda_{\rm st}$ as illustrated in Fig. 1. Thus, the LGI is violated due to Josephson oscillations when $\Lambda\neq 0$, whereas at $\Lambda=0$, the violation is due to single-particle oscillations. The violation of the LGI for $\Lambda\neq 0$ implies that macrorealism, which consists of MRPS and NIM, is excluded even for such a large $N$, and thus contradicts what is expected by the GP description; the LGI is never broken for the GP description [31]. In Figs. 1 and 1, the LGI is also violated at smaller values of $N$ for $|\Lambda|$ below the critical values near 2. These critical values coincide with the thresholds for the onset of self-trapping for respective $N$.
The self-trapping prevents the system from coherent oscillation and yields $C_{12}=C_{13}=1$ for any interval $\tau$ (see SM). Then $C_{12}$ and $C_{13}$ cancel out and $K=C_{23}\leq 1$. Thus, the LGI is nonviolated when self-trapping occurs for $|\Lambda|\geq\Lambda_{\rm c}(N)$.
Figure 1 shows the maximal violations of the LGI for particle numbers corresponding to Figs. 11. It indicates that the LGI is no longer violated, regardless of $\tau$, once $\Lambda$ exceeds the critical value $\Lambda_{\rm c}(N)$ for each $N$. Notably, this critical value approaches, but remains below, the threshold for the onset of self-trapping $\Lambda_{\rm st}^{\rm GP}=2$ as $N$ increases. Remarkably, the violation of the LGI becomes more significant as $N$ increases for fixed $\Lambda$, as shown in Figs. 1 and 1 (also see SM).
Zárate et al. investigated the violation of the LGI for two weakly coupled BECs in a double-well potential and found a maximal violation of the LGI 1.5 [15]. This is attributed to their presumption of the two-state approximation that results in two-outcome measurements [9] (also see SM). In contrast, Fig. 1 shows an LG violation exceeding 2, consistent with the result in Ref. [10], due to the use of multiple projectors |$\ket{n_{\rm L}}\ket{n_{\rm R}}$ and coarse graining of the results by the sign ($\hat{S}_{z}$).
Figures 1 and 1 show that the period of oscillation of $K$ increases and its amplitude decreases as $\Lambda$ approaches $\Lambda_{\rm c}(N)$ from below. It is due to the softening of the Josephson oscillation associated with the transition from the Josephson regime to the self-trapping regime [40, 46].
The LGI becomes nonviolated for $\tau$ greater than a characteristic value when $0<\Lambda<\Lambda_{\rm c}(N)$. This characteristic value of $\tau$ becomes smaller as $\Lambda$ increases for fixed $N$ and becomes larger as $N$ increases for fixed $\Lambda$, as shown in Figs. 11. This behavior is due to the damping of Josephson oscillation, which destroys the correlation in time. The Josephson oscillation indeed decays after a characteristic time, which becomes smaller as the interaction strength increases, as shown in Figs. 11, but it becomes longer as $N$ increases due to suppression of the fluctuation of the phase difference of two BECs.
The LGI is found to be violated symmetrically between repulsive ($\Lambda>0$) and attractive ($\Lambda<0$) on-site interaction in Figs. 11 since the Hamiltonian with a repulsive interaction can be mapped to the one with an attractive interaction as $\hat{R}_{z}(\pi)\hat{H}_{\rm BJJ}(\Lambda)\hat{R}_{z}^{\dagger}(\pi)=-\hat{H}_{\rm BJJ}(-\Lambda)$, where $\hat{R}_{z}(\theta)=\exp\quantity(-i\theta\hat{S}_{z})$. From this relationship, we obtain
$P(n_{\rm L},\Lambda,t)=P(n_{\rm L},-\Lambda,t),$ (3)
where $P(n_{\rm L},\Lambda,t)$ denotes the probability of $n_{\rm L}$ bosons on the left-well at time $t$ for $\Lambda$ (see SM). This symmetric relationship between repulsive and attractive on-site interaction results in the symmetric evolution of $K$ between them in Figs. 11.
Effect of the projective measurements on the LGI. The violations of the LGI in Figs. 11 arise from the absence of either MRPS, NIM, or both. Although it is not possible to conclusively determine whether the violation is specifically caused by one over the other, the violation of the LGI necessarily implies that at least one of these assumptions does not hold. Here, we examine the impact of projective measurements to investigate whether the LGI violation indicates a breakdown of NIM.
[…]
Figure 2: Effect of the projective measurements at time $t_{2}$. (a) Plot of $K$ as a function of $\tau J$. The upper blue region represents a violation of the LGI in Eq. (2). (b) $|\langle Q_{3}\rangle_{{\rm M}_{2}}-\langle Q_{3}\rangle|$ as a function of $\tau J$. The blue region indicates $\tau J$ at which the LGI is violated in (a). The parameters used in (a) and (b) are $N=100$ and $\Lambda=1$.
The effect of projective measurements appears explicitly in Fig. 2 in the comparison between $\langle Q_{3}\rangle_{{\rm M}_{2}}$, where the state is assumed to be collapsed due to measurements at $t_{2}$, and $\langle Q_{3}\rangle$, where measurements are not made at $t_{2}$. The condition $|\langle Q_{3}\rangle_{{\rm M}_{2}}-\langle Q_{3}\rangle|\neq 0$ implies a breakdown of NIM, indicating that the dynamics after time $t_{2}$ evolves differently depending on whether a measurement is made at $t_{2}$ or not. A comparison between Figs. 2 and 2 shows that whenever the LGI is violated, $|\langle Q_{3}\rangle_{{\rm M}_{2}}-\langle Q_{3}\rangle|$ is nonzero. Moreover, the greater the violation of the LGI, the larger is the value of $|\langle Q_{3}\rangle_{{\rm M}_{2}}-\langle Q_{3}\rangle|$ (also see SM). Note that $|\langle Q_{3}\rangle_{{\rm M}_{2}}-\langle Q_{3}\rangle|\neq 0$ does not necessarily lead to violation of the LGI. In Fig. 2, the oscillation of $K$ decays, and the LGI becomes nonviolated. Accordingly, the effect of projective measurements disappears in Fig. 2.
Observability of violation of the LGI.
[…]
Figure 3: (a) Time evolutions of $P(n_{\rm L})$ for $z(0)=0.28$. (b) Violations of the LGI for $z(0)=0.2$ (the red line), 0.28 (the blue line), and 0.4 (the green line). These parameters are well below the critical value of macroscopic quantum self-trapping. In (a) and (b), the parameters are $N=1000$ and $\Lambda=15$.
We numerically test the LGI for nearly the same parameters for which Josephson oscillations were observed in Ref. [21]. Here, we set $\Lambda=15$ and $N=1000$. The LGI is shown to be violated for realistic parameters in Fig. 3, in which the evolution of $K$ is plotted as a function of $\tau$ for several initial population imbalances $z(0)=(n_{\rm L}(0)-n_{\rm R}(0))/N$. $z(0)$ needs to be smaller than the critical value for self-trapping $z_{\rm st}(0)\simeq 0.5$ [33]. $z(0)=0.28$ is expected to be more desirable for the demonstration of violation of the LGI. Figure 3 shows the time evolution of $P(n_{\rm L})$ for $z(0)=0.28$. It exhibits the tunneling of atoms between the two wells, which indeed leads to the violation of the LGI.
Single-particle regime. Next, we study violation of the LGI in the absence of an on-site interaction. In this case the Hamiltonian is simply given by $\hat{H}_{\rm BJJ}=-J\hat{S}_{x}$. For this model, the time-dependent wave function can be obtained analytically. The wave function at time $t$ evolved from the initial state $|N-l\rangle_{\rm L}|l\rangle_{\rm R}$ is given by (see SM)
|$\ket{\Psi(t)}=$ $\frac{1}{\sqrt{(N-l)!}}\frac{1}{\sqrt{l!}}\left(\hat{a}^{\dagger}_{\rm L}\cos\frac{J}{2}t-i\hat{a}^{\dagger}_{\rm R}\sin\frac{J}{2}t\right)^{N-l}$
×$\times\left(\hat{a}^{\dagger}_{\rm R}\cos\frac{J}{2}t-i\hat{a}^{\dagger}_{\rm L}\sin\frac{J}{2}t\right)^{l}|\mathrm{vac}\rangle.$ (4)
[…]
Figure 4: (a) Violation of the LGI. (b) Three correlation functions $C_{12}$ (the red dashed line), $C_{23}$ (the black solid line), and $C_{13}$ (the green dotted line) as functions of $\tau$. (c)/(d) Time evolution of $P(n_{\rm L})$ in the absence/presence of the measurement at $tJ/2\pi=0.2$ (represented by the gray dotted line). $n_{\rm L}=70$ at $tJ/2\pi=0.2$ in (d). (e) Color map of $K$ for all different initial occupation numbers in the left well $n_{\rm L}(0)$. The violation of the LGI is symmetric under the reflection with respect to the line $n_{\rm L}(0)=50$. The blue color in (e) shows the part where the LGI is violated. In (a)–(e), the parameters are $N=100$ and $\Lambda=0$.
The first (second) line of Eq. (Violation of the Leggett-Garg inequality for dynamics of a Bose-Einstein condensate in a double-well potential) represents the coherent oscillation of the bosons initially located in the left (right) well. These bosons flow in opposite directions. The former represents bosons initially tunneling from the left well to the right well, as shown in Fig. 4, while the latter represents those initially tunneling from the right well to the left well. As shown in Fig. 4, the peak of the occupation probability coherently oscillates between $n_{\rm L}=0$ and $n_{\rm L}=N$ without damping. Due to this coherent oscillation the LGI is violated, as shown in Fig. 4.
The violation of the LGI in Fig. 4 closely resembles the one presented in Fig. 2(a) in Ref. [12]. This is due to the special coarse graining of the results as the sign ($S_{z}$). Our work represents a concrete realization of the proposal of Lambert et al. for the coarse-grained implementation of LGI violations.
Focusing on the first violation in Fig. 4, this behavior can be explained as follows: The LGI holds until $C_{13}$ starts to decrease, since $C_{12}=C_{13}=C_{23}=1$. The LGI begins to be first violated when $C_{13}$ starts decreasing at $\tau J/2\pi\simeq 0.125$. Since $C_{13}$ abruptly becomes $-1$ while $C_{12}=1$ and $C_{23}$ is still positive, the LGI becomes violated quite abruptly. The value of the peak of $K$ becomes larger with the increase of $N$, due to a more abrupt decrease of $C_{13}$ (see SM).
The persistence of this violation is attributed to the bosons tunneling backwards in Fig. 4 induced by the measurement, which corresponds to the second line of Eq. (Violation of the Leggett-Garg inequality for dynamics of a Bose-Einstein condensate in a double-well potential). In the case of $C_{23}$ the bosons tunneling backwards impede the decrease of $C_{23}$, while such an effect is absent in the case of $C_{13}$, where no measurement is performed at $t_{2}$ [see Fig. 4]. The LGI again begins to hold approximately when $C_{12}$ becomes zero at $\tau J/2\pi\simeq 0.25$ [see Fig. 4].
We also test the LGI for all different initial occupation numbers of the left well $n_{\rm L}(0)$ with $N=100$. In Fig. 4, it can be seen that the peak value of $K$ increases with a larger initial population imbalance (see SM).
Experimental realization. The experimental setup in Ref. [21] provides a concrete realization of the theoretical predictions in our paper. Phase contrast imaging (PCI) is a promising candidate for implementing our measurement scheme as a nondestructive measurement of a BEC in a double-well potential [47, 48, 49, 50]. PCI can perform measurements in the number basis |$\ket{n_{\rm L}}\ket{n_{\rm R}}$, enabling the measurement of the population in each well nondestructively without atom loss. This approach allows for the acquisition of each correlation function.
Conclusion and discussion. In this Letter, we study violations of the LGI for bosons in a double-well potential. We find that the LGI is violated due to Josephson oscillations in the weak-coupling regime, while it remains unviolated in the strong-coupling regime for $\Lambda$ beyond the critical $\Lambda_{\rm c}$, which depends on the particle number $N$. $\Lambda_{\rm c}$ approaches, but remains below, the critical threshold for the onset of self-trapping, $\Lambda_{\rm st}^{\rm GP}=2$, as derived from semiclassical analysis. It indicates that the nonviolation of the LGI is attributed to the self-trapping phenomena. Remarkably, we observe that the violation of the LGI becomes increasingly significant as the particle number increases. We also find that, in this system, an on-site interaction induces a dephasing effect that destroys temporal correlations, resulting in the LGI no longer being violated afterwards. Furthermore, we have revealed that the LGI is violated for nearly the same parameters for which the Josephson oscillations are observed in Ref. [21], thus indicating the presence of macroscopic quantum coherence in the semiclassical regime.
Acknowledgments. We thank anonymous reviewers for insightful comments. We thank G. Kimura, M. Kunimi, and T. Nikuni for fruitful discussions. S.T. thanks F. Dalfovo, G. Ferrari, and A. Recati for useful discussions. R.Y. is sup- ported by the Japanese Society for the Promotion of Science Grant-in-Aid for Scientific Research (KAKENHI, Grants No. 19K14616 and No. 20H01838) and of the WPI program “Sustainability with Knotted Chiral Meta Matter (SKCM2)” at Hiroshima University. S.T. is supported by the Japan Society for the Promotion of Science Grant-in-Aid for Scientific Research (KAKENHI Grant No. 19K03691).